The perimeter of a rectangle is 68 m. The base is six less than three times the height. What is the area?
Answer by rfer(16322)

2x+2(3x-6)=68
2x+6x-12=68
8x=80
x=10
3x-6=24
--------------
A=lw
A=24*10
A=240 sq m